OpenStudy (anonymous):

A and B are two bodies each with mass A kg and B kg. Initially the bodies are at rest. Equal forces are exerted to A and B for the same duration. The ratio between the final kinetic energy of A and B is : A. 1 B. B/A C. A/B D. (B/A)^2 E. (A/B)^2

OpenStudy (sidsiddhartha):

kinetic energy of a particle with mass m and velocity v is \[K_E=\frac{ 1 }{ 2 }mv^2\] so kinetic energy is directly proportional to the mass of the particle \[(K_E)_A/(K_E)B=\frac{ M_A }{M_B }=A/B\]

OpenStudy (anonymous):

thanks genuis !

OpenStudy (sidsiddhartha):

no prob :)